# NavRL: Learning Safe Flight in Dynamic Environments

### v1

@@ IV-A RL Training Results
Our framework utilizes a curriculum learning strategy for our RL policy training, as illustrated by the initial and final training environments shown in Fig. 6. Static obstacles are depicted in red and dynamic obstacles are shown in green. Based on our observations, dynamic obstacles present a greater challenge during training than static obstacles. Therefore, we gradually increase the number of dynamic obstacles in the environment from 60 to 120, in increments of 20, once the navigation success rate exceeds 80% (defined as safely navigating from the start to the goal without collisions). To demonstrate the effectiveness of curriculum learning, Table I compares the highest navigation success rates achieved with and without curriculum learning, keeping the total training time constant for both scenarios. The table shows that as the number of dynamic obstacles increases, the highest navigation success rate without curriculum learning decreases more sharply compared to training with curriculum learning, highlighting its effectiveness. We stopped policy training at 120 dynamic obstacles and saved the best model trained with 100 dynamic obstacles, as it achieved a relatively high success rate (80.96%) under challenging conditions.
@@
To show the importance of training with a larger number of robots, Fig. 8 compares the RL training returns achieved with different numbers of robots deployed in training. The figure shows that training with more robots not only leads to faster convergence speed but also results in higher RL returns. Our experiments were conducted with 1024 robots, which utilized the maximum available GPU memory.
@@
Figure 8: Plot of the smoothed training return curves. Training with a larger number of robots leads to faster convergence and achieves a higher return.
@@ IV-B Simulation Experiments
We designed several simulation experiments to evaluate the performance of the proposed framework. All simulation experiments were conducted in Gazebo to demonstrate the framework’s sim-to-sim transferability. We first conducted quantitative experiments in multiple indoor-like dynamic environments to simulate scenarios where the robot must navigate safely around humans. These environments are intended to resemble real-world scenarios and are less complex than the training environments. Fig. 7 illustrates a safe trajectory resulting from our framework in one of the example environments. From all the experiments, we observed that the robot successfully performed navigation without collisions.
To quantitatively evaluate the performance, we generate environments with high obstacle density, mirroring the conditions of our training environments, where both static and dynamic obstacles are randomly placed. Given the limited availability of open-source RL-based navigation benchmark algorithms, we compare our method with the popular optimization-based static planner [54] and a vision-aided planner [20] designed for dynamic environments. Additionally, we evaluate the performance of our framework with and without the safety shield to verify its effectiveness. Specifically, we test in three types of environments: static, purely dynamic, and hybrid environments. Each algorithm is run 20 times in each environment, and the average number of collisions per run is calculated as shown in Table II. It is important to note that these environments are designed to test the limits of the algorithms and are significantly more complex than real-world scenarios. Therefore, the occurrence of collisions does not necessarily indicate that the algorithm is unsafe. Overall, our NavRL demonstrates the lowest number of collisions in dynamic and hybrid environments, while maintaining a comparable collision rate to the EGO planner in static environments. The table shows N/A for the EGO planner in dynamic and hybrid environments because its inefficient map updates cause it to get stuck when handling excessive noise from dynamic obstacles. In the experiments, ViGO fails to provide sufficiently reactive trajectories for collision avoidance when obstacles are in close proximity to the robot. In contrast, our method can generate more reactive controls that help the robot avoid collisions more efficiently. Comparing our framework with and without the safety shield, we find that the safety shield reduces the number of collisions across all environments and performs particularly well in purely dynamic settings. Since avoiding dynamic obstacles is more difficult, making the network more prone to failure and mistakes, the safety shield effectively prevents these errors.
@@
Avg. Collision Times Measurement in a 20m×$\times$40m Map
@@
Ours (NavRL) 0.65 ($81.3\%$) 0.85 ($27.0\%$) 2.10 ($47.8\%$)
@@ IV-C Physical Flight Tests
To demonstrate the sim-to-real transfer capability and safe navigation performance in the real world, we conducted several physical flight experiments in various settings, as shown in Fig. 9. In these experiments, static obstacles were placed in the environments, and several pedestrians were directed to walk toward the robot, requiring the robot to actively avoid collisions while achieving its goals. Observations from the flight experiments demonstrate that the robot successfully avoided collisions and reached its goal locations safely.
@@ V Conclusion and Future Work
This paper presents a novel deep reinforcement learning framework, NavRL, designed to achieve safe flight in dynamic environments based on the Proximal Policy Optimization (PPO) algorithm. The framework features specially designed state and action representations that enable the learned policy to make safe navigation decisions in the presence of both static and dynamic obstacles, facilitating effective zero-shot transfer from simulation to real-world scenarios. Besides, the proposed framework incorporates a safety shield based on the velocity obstacle concept to avoid severe failures caused by the black-box nature of neural networks. Moreover, by employing a parallel training pipeline using NVIDIA Isaac Sim, the framework significantly accelerates the learning process. Results from simulation and physical experiments verify the effectiveness of our approach in achieving safe navigation within dynamic environments. Future work will focus on improving and adapting this framework for deployment across various robotic platforms.

### v2

@@ IV-A RL Training Results
Our framework utilizes a curriculum learning strategy for our RL policy training, as illustrated by the initial and final training environments shown in Fig. 6. Static obstacles are shown in red with a color gradient indicating height, while dynamic obstacles are depicted in green. Based on our observations, dynamic obstacles present a greater challenge during training than static obstacles. Therefore, we gradually increase the number of dynamic obstacles in the environment from 60 to 120, in increments of 20, once the navigation success rate exceeds 80% (defined as safely navigating from the start to the goal without collisions). To demonstrate the effectiveness of curriculum learning, Table I compares the highest navigation success rates achieved with and without curriculum learning, keeping the total training time constant for both scenarios. The table shows that as the number of dynamic obstacles increases, the highest navigation success rate without curriculum learning decreases more sharply compared to training with curriculum learning, highlighting its effectiveness. We stopped policy training at 120 dynamic obstacles and saved the best model trained with 100 dynamic obstacles, as it achieved a relatively high success rate (80.96%) under challenging conditions.
@@
To show the importance of training with a larger number of robots, Fig. 8 compares the average RL training returns achieved with different numbers of robots deployed in training. The figure shows that training with more robots not only leads to faster convergence but also results in higher RL returns. Our experiments were conducted with 1024 robots, which utilized the maximum available GPU memory.
@@
Figure 8: Plot of the smoothed average training return curves. Training with a larger number of robots leads to faster convergence and higher returns.
@@ IV-B Simulation Experiments
We conducted simulation experiments in Gazebo to evaluate the proposed framework and demonstrate its sim-to-sim transferability. Quantitative tests were performed in indoor-like dynamic environments resembling real-world scenarios where the robot navigates safely around humans. Fig. 7 shows an example of a safe trajectory. The results confirm that the robot successfully navigated without collisions.
To quantitatively evaluate the performance, we generate environments with high obstacle density, mirroring the conditions of our training environments, where both static and dynamic obstacles are randomly placed. Given the limited availability of open-source RL-based navigation benchmark algorithms, we compare our method with the popular optimization-based static planner [48] and a vision-aided planner [20] designed for dynamic environments. Additionally, we evaluate the performance of our framework with and without the safety shield to verify its effectiveness. Specifically, we test in three types of environments: static, purely dynamic, and hybrid environments. Each algorithm is run 20 times in each environment, and the average number of collisions per run is calculated as shown in Table II. The percentage value in Table II indicates the average collision times relative to the baseline method [20]. It is important to note that these environments are designed to test the limits of the algorithms and are significantly more complex than real-world scenarios. Therefore, the occurrence of collisions does not necessarily indicate that the algorithm is unsafe. Overall, our NavRL demonstrates the lowest number of collisions in dynamic and hybrid environments, while maintaining a comparable collision rate to the EGO planner in static environments. The table shows N/A for the EGO planner in dynamic and hybrid environments because its inefficient map updates cause it to get stuck when handling excessive noise from dynamic obstacles. In the experiments, ViGO fails to provide sufficiently reactive trajectories for collision avoidance when obstacles are in close proximity to the robot. In contrast, our method can generate more reactive controls that help the robot avoid collisions more efficiently. Comparing our framework with and without the safety shield, we found the safety shield consistently reduced collisions, particularly in dynamic environments, where it mitigates errors caused by the neural network’s increased failure risk.
@@
Average Collision Times Measurement in a 20m×$\times$40m Map
@@
Ours (NavRL) 0.65 ($\mathbf{81.3\%}$) 0.85 ($\mathbf{27.0\%}$) 2.10 ($\mathbf{47.8\%}$)
@@ IV-C Physical Flight Tests
To demonstrate sim-to-real transfer and safe navigation, we conducted physical flight experiments in various settings, as shown in Fig. 9. Static obstacles were placed in the environments, and several pedestrians were directed to walk toward the robot, requiring it to avoid collisions while reaching its goals. The results showed that the robot successfully avoided collisions and reached its destinations safely.
@@ V Conclusion and Future Work
This paper presents a novel deep reinforcement learning framework, NavRL, designed to achieve safe flight in dynamic environments based on the Proximal Policy Optimization (PPO) algorithm. The framework uses tailored state and action representations to enable safe navigation in both static and dynamic environments, supporting effective zero-shot sim-to-sim and sim-to-real transfer. Besides, A safety shield based on the velocity obstacle concept mitigates failures from the black-box nature of neural networks. Additionally, a parallel training pipeline with NVIDIA Isaac Sim accelerates the learning process. Results from simulation and physical experiments verify the effectiveness of our approach in achieving safe navigation within dynamic environments. Future work will focus on improving and adapting this framework for deployment across various robotic platforms.